# Emergent photons and mechanisms of confinement

### v1

@@ III Numerical simulations and results
They argued that for a particular choice of $\beta$ and $\tilde{\beta}$, a massless photon phase emerges, but their numerics were inconclusive. We computed the smeared Wilson loops and present the results in Fig. 4. We computed the smeared Wilson loops in a 3x8 grid of values $\tilde{\beta}\times\beta\in[-0.379,0.407]\times[1.4,1.6]$ by running $10^{4}$ Monte Carlo sweeps of an $L=6$ lattice. We present the time series histogram of the smeared Wilson loops in Fig. 4. As can be observed, we see the same pattern of circle formation indicating a photon phase in between the confined and the SSB phase, confirming the finding of [22].
Figure 4: The smeared Wilson loop as a function of $\beta$ and $\tilde{\beta}$ (see (3)) of the $\mathbb{Z}_{4}$ lattice gauge theory.
Lastly, we take a look at the $\mathbb{Z}_{3}$ model. In this model with the action (1), there are only two phases. The picture of this model is that as $\beta$ decreases from $\beta\gg 1$ to $\beta\ll 1$, both center vortices and monopole-junctions become abundant simultaneously. To open a photon phase, we therefore suppress the monopoles and write the action
$S=-\beta\sum_{p}\text{Re\;}W_{p}-\mu\sum_{c}(dm)_{c}^{2}\;.$ (4)
To identify a third phase, we look at the action susceptibility (i.e. heat capacity) defined as
In Fig. 3 we plot numerical Monte Carlo results $\chi_{S}$ as a function of $\beta$ at $\mu=1$. We ran Monte Carlo simulations for $L=8,10,12,$ and $14$ for various values of $\beta$, sampling more densely around the two suspected critical values. Moreover, a monopole suppression of $\mu=1$ was chosen. For each experiment, $4\cdot 10^{4}$ measurements were performed, with 100 decorrelation sweeps between them. The error was estimated using Bootstrap. The results for $\chi_{S}$ can be seen in Fig. 3. As is clear, the susceptibility shows two peaks, indicating a three-phase scenario, with phase transitions at $\beta=\beta_{1}\approx 0.265$ and $\beta=\beta_{2}\approx 0.512$. The inset shows the smeared Polyakov loop $\bar{P}$ for values of beta close to the second transition, showing the circles forming in the photon phase.
Finally, we also consider the connected plaquette-plaquette correlator (PPC) defined as
$C_{x-y;\mu\nu}=\left\langle W_{p_{x,\mu\nu}}^{*}W_{p_{y,\mu\nu}}\right\rangle-\big{\langle}W_{p_{x,\mu\nu}}^{*}\big{\rangle}\big{\langle}W_{p_{y,\mu\nu}}\big{\rangle}\;,$ (6)
where $p_{x,\mu\nu}$ indicates a plaquette with a corner at lattice site $x$, expanding in directions $\mu,\nu=0,1,2,3$, and averages are over ensemble and all spatial translations. At long distances in the photon phase, the theory flows to a pure $\text{U}(1)$ gauge theory with the action $S=\frac{1}{4}\int d^{4}x\;F_{\mu\nu}F^{\mu\nu}$, which describes free photons. All gauge invariant local operators of the $\text{U}(1)$ gauge theory must be built from the field-strength operator $F_{\mu\nu}$, so we expect $C_{x-y;\mu\nu}\propto\left\langle F_{\mu\nu}(x)F_{\mu\nu}(y)\right\rangle$ for large $|x-y|$. Further, we have that
where $D_{\mu\nu;\rho\sigma}$ is independent of the modulus $|x-y|$ and is easily computed [^4] One way is to note that $A_{\mu}(x)A_{\nu}(y)=\delta_{\mu\nu}/|x-y|^{2}$ in the Feynman gauge, and then take derivatives., but is not relevant for our discussion. Note that if points $x$ and $y$ differ by translation along the time direction $x^{0}$, the correlator $\left\langle F_{ij}(x)F_{ij}(y)\right\rangle>0$, while $\left\langle F_{0i}(x)F_{0i}(y)\right\rangle<0$ by reflection positivity, so it makes sense to look at the positive definite combination of the two, which on the lattice corresponds to
$C(n)=\sum_{\rho}\Big{(}\sum_{\begin{subarray}{c}\mu>\nu\\ \mu,\nu\neq\rho\end{subarray}}C_{n\hat{\rho},\mu\nu}-\sum_{\mu\neq\rho}C_{n\hat{\rho},\rho\mu}\Big{)}\approx\frac{K}{n^{4}}\;,$ (8)
valid up to corrections $\mathcal{O}(1/n^{6})$, where $\hat{\rho}$ is a lattice unit vector in the direction $\rho$, and $K$ is a constant. We ran Monte Carlo simulations of $\mathbb{Z}_{3}$ LGT with $\mu=1$, $\beta=0.5114$ and $L=16$, performing $1.28\cdot 10^{6}$ measurements, with 500 decorrelation sweeps between them. The error was estimated using a binning Jackknife analysis. The quotient $\frac{C(n)}{C(n+1)}$, behaves as $\frac{(n+1)^{4}}{n^{4}}+o(1/n^{2})$ for large $n$ and has no fitting parameter. We compare this theoretical curve with lattice results in Fig. 5 and find excellent agreement, confirming the photon phase of the model.
Figure 5: Lattice data shows the ratio of correlators $C(n)/C(n+1)$ and the smooth red curve shows the theoretical prediction assuming 4th power law decay, consistent with the occurrence of photon phase. Note that the curve has no fitting parameters.
IV Conclusions
We numerically showed that $\mathbb{Z}_{N}$ lattice gauge theories have three phases for all $N\geq 3$, supplementing the spontaneously broken and restored $\mathbb{Z}_{N}$ 1-form symmetry with a phase of emergent photons, where the symmetry enhances to $U(1)$.
In addition, we demonstrated a direct link between the photon phase and the proliferation of string-like center vortices, while the confined phase was linked to proliferation of both vortices and vortex junctions (monopoles) as proposed in [7]. This suggests that center vortices are not directly responsible for confinement as commonly believed for $SU(N)$ gauge theory. [^5] Defining monopoles and center vortices requires gauge fixing in $SU(N)$ gauge theories (see e.g. [23, 10, 11]), although there are some efforts to identify such objects without it [24, 24]. Further, a novel proposal of how to discretize $\text{SU}(N)$ gauge theories [25, 26], may be related to a gauge invariant definition of vortices and monopoles [27].
The case of $\mathbb{Z}_{2}$ lattice gauge theory is more subtle as it does not have a global charge conjugation symmetry $C$ [^6] TS would like to thank M. Nguyen and E. Richards for discussions on this point. While it was proposed in [7] to have a photon phase when deformed, such deformations would introduce the $C$-symmetry. A similar issue occurs in $D_{N}$ lattice gauge theories, which have no $C$ symmetry and only at most a $\mathbb{Z}_{2}$ 1-form symmetry. Yet for large enough $N$, $D_{N}$ LGT should be indistinguishable from $O(2)$ LGT, which certainly has a photon phase, albeit with a non-invertible emergent 1-form symmetry. The Landau paradigm for $\mathbb{Z}_{2}$ symmetry therefore likely involves a massless phase as suggested in [7], although the details are more subtle and are left for future explorations.
Acknowledgement
We would like to thank Mathew Bullimore, Xavier Crean, Christof Gattringer, Tyler Helmuth, Nabil Iqbal, Biagio Lucini, Mendel Nguyen, Emily Richards, Shu-Heng Shao and Yuya Tanizaki for the discussions and comments. This work has made use of the Hamilton HPC Service of Durham University. TS is supported by the Royal Society University Research Fellowship and, in part, by the STFC consolidated grant number ST/T000708/1. JG is supported by EPSRC grant EP/Y028872/1.

### v2

@@ Appendix C Smeared Polyakov Loops
Figure 7: Illustration of the transformation $s_{\ell}\to s_{\ell}+1$ on the links crossing the hyperplane $H$. The left panel shows the original configuration, and the right panel highlights the links affected by the shift.
This observation justifies augmenting our dataset for the smeared Polyakov histograms by including the $N$ copies of $\overline{P}$ obtained through rotations by angles of $2\pi m/N$ with $m=0,\dots,N-1$.
Appendix D Reducing Floating-Point Error
One of the big advantages of working with $\mathbb{Z}_{N}$ is that many observables can only take on a finite set of values, often of the size of $N$. This discreteness can be exploited to store measurements compactly as integer counts, avoiding cumulative sums of floating points and the rounding errors they introduce. For example, the non-connected piece in the correlator includes a term $W^{*}_{p_{x,\mu\nu}}W_{p_{y,\mu\nu}}$ which reduces to averages of phases of the form $e^{i\frac{2\pi(n_{x}-n_{y})}{N}}$. As the differences $n_{x}-n_{y}$ only take values in $\mathbb{Z}_{N}$, and carry all the information, storing these values encodes the data concisely and without loss of precision. We applied the same strategy to record Polyakov loops and average plaquette measurements. In the final data analysis step, the integer arrays can then be processed at the desired precision and turned into concrete measurements.
Appendix E Correlators
In this section, we present a more detailed analysis of the correlators. We defined two plaquette-plaquette correlators, which we repeat here for convenience
$C^{\pm}_{x-y;\mu\nu}=\left\langle W_{p_{x,\mu\nu}}^{\pm 1}W_{p_{y,\mu\nu}}\right\rangle-\big\langle W_{p_{x,\mu\nu}}^{\pm 1}\big\rangle\big\langle W_{p_{y,\mu\nu}}\big\rangle\;,$ (9)
where there is no sum over the repeated indices $\mu,\nu$ on the RHS. We wish to sum over indices $\mu$ and $\nu$, but here we need to distinguish between the longitudinal and transverse correlators, defining them as follows
$C^{\pm}_{L}(n)=\frac{1}{4}\cdot\frac{1}{3}\sum_{\begin{subarray}{c}\rho,\mu\\ \rho\neq\mu\end{subarray}}C^{\pm}_{n\hat{\rho},\rho\mu}$ (10a)
$C^{\pm}_{T}(n)=\frac{1}{4}\cdot\frac{1}{3}\sum_{\begin{subarray}{c}\rho,\mu,\nu\\ \mu>\nu\neq\rho\end{subarray}}C^{\pm}_{n\hat{\rho},\mu\nu}$ (10b)
where $\hat{\rho}$ is the unit vector in the direction of $\rho$. The factors of $\frac{1}{4}\cdot\frac{1}{3}=\frac{1}{12}$ are to account for the averages over indices.
Note that by reflection-positivity [32] we expect that $C_{L}^{+}(n)$ and $C_{T}^{-}(n)$ are each positive, so it makes sense to add them up for increased statistics. We hence define the total correlator to be
$C(n)=C_{L}^{+}(n)+C_{T}^{-}(n)\;.$ (11)
Up to the factor of $1/12$, this is the same correlator defined in the main article used in the ratio $C(n+1)/C(n)$.
In the main text, we presented the ratios between the correlators of subsequent lattice spacing. This analysis revealed that the fourth-order contributions dominate the long-range scaling behavior. Nevertheless, we wish to show that the curves are fitted well with the power law ansatz
$\sum_{k=2}^{\infty}a_{2k}\!\left(\frac{1}{x^{2k}}+\frac{1}{(L-x)^{2k}}\right),$
In Fig. 8 we show the measurements of correlators $C_{L}^{+},C_{T}^{-}$ and $C$, as well as a fitted curve up to the order $1/x^{10}$ for the total correlator. The fitting parameters are shown in Table 2, where we skipped the 2 smallest $n$ correlator data points.
The data is the same as the one used for Fig. 5, i.e. we ran Monte Carlo simulations for $\mathbb{Z}_{3}$ with $\beta=0.5114$ and $\mu=1$, for $\mathbb{Z}_{4}$ with $\beta=1.514$ and $\tilde{\beta}=-0.393$, and $\mathbb{Z}_{7}$ with $\beta=1.8$. A toric lattice of size $L=16$ was used, and we performed $1.28\cdot 10^{6}$ measurements, with 500 decorrelation sweeps between them. The error was estimated using a Bootstrap.
Raw correlators[Uncaptioned image][Uncaptioned image][Uncaptioned image]
Figure 8: Raw correlators for the longitudinal, transverse, and total contributions, obtained from Monte Carlo simulations on a toric lattice of size $L=16$. Results are shown for $\mathbb{Z}_{3}$ at $\beta=0.5114$ with $\mu=1$, for $\mathbb{Z}_{4}$ at $\beta=1.514$ with $\tilde{\beta}=-0.393$, and for $\mathbb{Z}_{7}$ at $\beta=1.8$. The curve for the total correlator is obtained by fitting the series expansion up to order 10.
NN $a_{4}$ $a_{6}$ $a_{8}$ $a_{10}$
$\mathbb{Z}_{3}$ 0.0913 0.4163 6.0466 -42.3533
$\mathbb{Z}_{4}$ 0.1023 0.4600 11.1962 -74.4772
$\mathbb{Z}_{7}$ 0.0825 0.5547 7.1300 -55.5292
Table 2: Fitting coefficients $a_{4}$, $a_{6}$, $a_{8}$, $a_{10}$ for the total correlator.
Appendix F Some comments on the robustness of the Goldstone boson phase
In this section, we contrast the Goldstone phases of $1$-form and 0-form symmetries. The main points have been discussed in the literature (see e.g. [2, 5, 33, 34, 35]), but we review them here for completeness.
We begin by briefly reviewing the Goldstone boson phase for ordinary, 0-form symmetries. Consider a system with an exact ordinary (i.e. 0-form) $\text{U}(1)$ symmetry. Further, let this system be pushed into a spontaneously broken phase, i.e. the Goldstone boson phase. The field $\varphi(x)$ corresponds to a Goldstone boson particle, i.e. an angle-valued field $\varphi\sim\varphi+2\pi$ which shifts under the exact $\text{U}(1)$ symmetry as follows
$\text{U}(1):\;\;\varphi(x)\rightarrow\varphi(x)+\alpha\;.$ (12)
The symmetry transformation parameter $\alpha$ is also angle-valued, i.e. $\alpha\sim\alpha+2\pi$, so the symmetry is a $\text{U}(1)$ symmetry.
Since we require the symmetry to be exact, the low-energy effective theory of the Goldstone boson must be at most derivatively coupled. Assuming full Lorentz symmetry, the effective Lagrangian is given by
$\mathcal{L}_{\mathit{eff}}=\frac{f}{2}(\partial_{\mu}\varphi)^{2}+(\text{Higher derivative terms})\;.$ (13)
Since higher derivative terms are irrelevant, the effective theory is that of a free compact scalar. We should ask now: is the massless Goldstone phase stable? Firstly we need to notice that the theory actually enjoys a (d2)(d-2)-form symmetry following conservation of the current $j_{\mu_{1}\mu_{2}\cdots\mu_{d-1}}=\frac{1}{2\pi}\epsilon_{\mu_{1}\mu_{2}\cdots\mu_{d-1}\mu_{d}}\partial^{\mu_{d}}\varphi(x)$. This symmetry should be considered as an emergent one, and it is explicitly broken by vortex defects. Importantly, in space-time dimension [^8] In $d=2$ there is, of course, no strict spontaneous symmetry breaking of continuous symmetries due to the Mermin-Wagner theorem, but there still can be massless particles. $d=2$, this is an ordinary 0-form symmetry, and vortex operators could drive the system to a gapped phase. In higher dimensions, vortices are extended objects, and their proliferation signals a phase transition to a symmetry-restored phase.
Let us therefore assume that the vortices are heavy or irrelevant, and that the $\text{U}(1)$ symmetry is explicitly broken to $\mathbb{Z}_{N}$. Could a Goldstone phase emerge? To answer this question, we must check whether the Goldstone phase has relevant perturbations which reduce the symmetry down to $\mathbb{Z}_{N}$. The simplest such perturbation is by a $\cos(N\varphi)$ operator. But such deformations are relevant whenever space-time dimension d>2d>2999In dimension $d=2$ the relevance of the operator depends on $f$ and its (ir)relevance is the driving mechanism of the Berezinskii-Kosterlitz-Thouless transition, and Goldstone bosons are lifted.
Now we repeat the argument for $\text{U}(1)$ 11-form symmetries. The Goldstone boson of such a symmetry is a $\text{U}(1)$ gauge field with an effective Lagrangian
$\mathcal{L}_{\mathit{eff}}=\frac{1}{4e^{2}}\mathcal{F}_{\mu\nu}^{2}+(\text{higher gauge-invariant derivatives)}\;,$ (14)
where $\mathcal{F}_{\mu\nu}=\partial_{\mu}\mathcal{A}_{\nu}-\partial_{\mu}\mathcal{A}_{\nu}$ is the curvature tensor [^10] Of course, $\mathcal{A}_{\mu}$ is not necessarily globally defined, and in general it must be defined in patches.. In other words, it is just a photon (or Coulomb) phase.
The $1$-form symmetry is a shift symmetry $\mathcal{A}_{\mu}(x)\rightarrow\mathcal{A}_{\mu}(x)+\alpha_{\mu}(x)$, where $\alpha_{\mu}(x)$ is a closed, but not exact $1$-form, i.e. a $1$-form with a vanishing curvature $\partial_{\mu}\alpha_{\nu}-\partial_{\nu}\alpha_{\mu}=0$. Note that $\alpha_{\mu}(x)$ is a member of the first cohomology group with coefficients in $\text{U}(1)$, exactly what we expect for the $1$-form symmetry. Furthermore, the Wilson loops transform with a $\text{U}(1)$ phase, i.e. $W[C]\rightarrow W[C]e^{i\int_{C}\alpha}$ – where $C$ is the space-time contour of the Wilson loop – implying that the symmetry is $\text{U}(1)$.
Before proceeding to break the $\text{U}(1)$ 1-form symmetry explicitly, we must ask whether the Goldstone boson theory (14) is robust. The theory enjoys the conservation of the multi-index current $j_{\mu_{1}\cdots\mu_{d-2}}=\frac{1}{4\pi}\epsilon_{\mu_{1}\cdots\mu_{d-1}\mu_{d}}F^{\mu_{d-1}\mu_{d}}$, which implies that the Goldstone boson theory (14) has an emergent (d3)(d-3)-form symmetry which is broken by extended (d3)(d-3)-dimensional defects called monopoles. Monopoles are analogous to vortices of theory (13). The monopoles are local operators in $d=3$ and, importantly, they are relevant, gapping out the Goldstone bosons. Therefore, the remainder of the discussion will then focus on $d=4$, where monopoles are extended, particle-like objects, which can be heavy, making the Goldstone boson phase robust.
As before, now we imagine that the $U(1)$ 11-form symmetry is not exact, but that it emerges, enhancing the $\mathbb{Z}_{N}$ 11-form symmetry. There is no operator we can add to (14) to achieve this, and the simplest thing we can do is add matter fields coupled to the $\mathcal{A}_{\mu}$ gauge field. Let us add a complex scalar field $\Phi(x)$, with a covariant derivative $D_{\mu}\Phi=(\partial_{\mu}+iN\mathcal{A}_{\mu})\Phi$, i.e. effective Lagrangian takes the form
$\mathcal{L}_{eff}=\frac{1}{4e^{2}}\mathcal{F}_{\mu\nu}^{2}+|D_{\mu}\Phi|^{2}+m^{2}|\Phi|^{2}+\cdots$ (15)
Note that the choice $\alpha_{\mu}=\frac{\partial_{\mu}\xi}{N}$, with $\xi(x)\sim\xi(x)+2\pi$ an angle-valued field, is still a symmetry of the resulting effective theory if we shift $\Phi\rightarrow\Phi e^{-i\xi(x)}$. This symmetry shifts the Wilson loop $W[C]$ by a $\mathbb{Z}_{N}$ phase and is hence a $1$-form $\mathbb{Z}_{N}$ symmetry. We therefore have achieved explicit breaking of the $\text{U}(1)$ 11-form symmetry down to the $\mathbb{Z}_{N}$ subgroup.
Crucially, as opposed to the 0-form symmetry Goldstone bosons, our attempt at breaking the $1$-form symmetry need not destroy the Goldstone boson phase. Indeed, if $m^{2}$ is large and positive, the matter field $\Phi$ can be integrated out, with minimal effect on the massless Goldstone bosons. This is why the Goldstone boson phase of the $1$-form symmetry is robust.
Appendix G Comment on the order of the phase transitions
The main text discusses the existence of three generic phases of a $\mathbb{Z}_{N}$ lattice gauge theory: the spontaneously broken phase (SSB), the restored (confined) phase, and the photon phase. While the nature of the phase transitions between these phases is not the main focus of the paper, this question is potentially an important one because, if they are continuous transitions, they could potentially furnish a UV completion of an interacting $\text{U}(1)$ gauge theory. We will, however, argue that we generically expect such transitions to be 1st order.
First, consider a transition from the SSB phase to the photon phase. Such a transition is in the same universality class as the condensation of a charge $N$ Higgs field of a $\text{U}(1)$ gauge theory. Such a transition was studied long ago by Coleman and Weinberg [36] (see also [37] for a more recent analysis) with the conclusion that it is always 1st order.
The photon/confinement transition is also the same type of transition (this time with a charge 1 Higgs field) by the electric-magnetic duality.
Further, in [6] it was argued that the Ginzburg-Landau theory of 1-form symmetries generically has a cubic term, and the authors believe this causes all transitions involving a 1-form symmetry to generically be 1st order. Further still, if the transition is 2nd (or higher) order, it must flow to a conformal fixed point in the IR. However, in [5] it was argued that a conformal field theory with a $\text{U}(1)$ 11-form symmetry (expected to emerge at the fixed point) must flow to a non-interacting photon phase, which has no relevant deformation driving it to a gapped phase.
A potential loophole is a transition where both electric and magnetic matter is light. This was the question asked in ref. [37], but no such fixed point was found with a single species of electric and magnetic matter. Nevertheless, the question remains an open one.
By the above arguments on fairly general grounds, we expect that all transitions observed in the $\mathbb{Z}_{N}$ lattice gauge theory are always 1st order, although we cannot preclude more exotic scenarios. The most interesting case to study would be at the point where all three phases meet, and where a strongly interacting fixed point cannot be a priori excluded.